# Part-Time Penalties and Heterogeneous Retirement Decisions

### v1

@@
(March 23, 2025)
Older male workers exhibit diverse retirement behaviors across occupations and respond differently to policy changes, influenced significantly by the part-time penalty—wage reduction faced by part-time workers compared to their full-time counterparts. Many older individuals reduce their working hours, and in occupations with high part-time penalties, they tend to retire earlier, as observed in data from Japan and the United States. This study develops a general equilibrium model that incorporates occupational choices, endogenous labor supply, highlighting that the impact on the retirement decision is amplified by the presence of assets and pensions. I find that cutting employees’ pension benefits reduce aggregate labor supply in occupations with high part-time penalties, reducing overall welfare across the economy. In contrast, increasing income tax credits and exempting pension form income tax boost labor supply across all occupations and enhance welfare by raising disposable wages relative to the reservation wage. Reducing part-time penalties in high-penalty occupations also stimulate the labor supply in high-penalty occupations and improve long-term welfare.
Placeholder
@@ 1 Introduction
Retirement timings vary significantly across occupations. While some professions tend to see workers retiring at sixty, others have mean retirement ages extending beyond seventy, as shown in Figure 1. Despite this striking heterogeneity, the factors driving such diverse retirement decisions remain unclear. Understanding these underlying factors is a requisite step to examine the impact of policy measures aimed at increasing the labor supply of older workers.
@@
This issue is particularly pressing as labor force is projected to decline further: working-age population in OECD countries is expected to shrink by 11% by 2062 compared to 2022 OECD [2023]. Given the prevalence of aging, which may amplify the variance in retirement decisions across occupations, it is plausible that policymakers should design retirement policies tailored to specific occupations.
This paper is the first to propose a framework to analyze how different occupations respond to policy changes, highlighting part-time penalties—wage reductions experienced by part-time workers compared to their full-time counterparts—as a key source of this heterogeneity. As illustrated in Figure 3, there is a negative correlation between part-time penalties and share of old workers: a proportion of male workers aged 60 and over among those aged 40 and over. The denominator includes both middle-aged and older workers to mitigate the influence of trends in occupational choice. Occupations with smaller part-time penalties tend to have higher rates of old workers as workers face less significant wage reductions when reducing working hours to spend more leisure time as they get older. In other words, in such occupations, workers are likely to retire earlier than in other occupations. This mechanism plays a crucial role in explaining the divergence in retirement timings across occupations.
@@
To classify occupations and compute moments for quantitative analysis, this study utilizes the Japanese Panel Study of Employment Dynamics (JPSED). The JPSED covers more than 200 occupations and provides detailed personal information on each worker, including birth year, sex, education, work history, family status, and more. These rich variables enable precise regressions for classifying occupations. Additionally, the Japanese Household Panel Survey (JHPS/KHPS) supplements the analysis with asset data, which is not available in the JPSED. To ensure that similar retirement behaviors are observed in other countries, the IPUMS-CPS (Flood et al. [2024]), which provides data for the United States, is also employed.
@@
I propose several unconventional policies to increase the labor supply of older workers while positively impacting overall welfare. For instance, reducing the part-time penalties in nonlinear occupations increases labor supply in the nonlinear occupations by $2.247\%$ and enhances long-term welfare in terms of consumption equivalence. Policies such as raising tax credits and exempting pensions from income tax are also effective in boosting labor supply across both types of occupations, increasing output, and improving welfare.
@@
Figure 3: Part-time Penalty and Share of Old Workers: Males, 2015-2019
@@ 2.1 Data
The Japanese Panel Study of Employment Dynamics (JPSED), collected by the Recruit Works Institute and released by the University of Tokyo, provides data for classifying occupations and calibrating the model. Covering over 100 individual attributes from 2015 to 2022, it includes approximately 50,000 annual observations. For classification, all valid observations are used with controls for time and sex. For calibration, the analysis focuses on males aged 25–104 from 2015 to 2019, avoiding the COVID-19 pandemic’s heterogeneous impact on retirement decisions.
@@
For robustness checks, IPUMS-CPS data from 2009 to 2024 is used to examine whether a similar pattern is observed in the U.S. The dataset, which includes approximately 130,000 to 220,000 individuals, provides detailed information on personal attributes and work-related characteristics. Occupations are classified based on this data, allowing for a comprehensive analysis of labor market trends and retirement behavior.
@@ 2.2 Occupational Classification Strategy
    For each of the more than 200 occupations, I regress hourly wages on a quartic polynomial of weekly working hours, controlling for age (as a polynomial), sex, education, marital status, child status, residential area, and time-fixed effects. Occupations with fewer than 200 observations are excluded.
@@
    I calculate the residualized hourly wage difference between individuals working 10 hours per week and those working 40 hours per week. I define this wage difference standarized by the average hourly wage of males aged 25-79 as a part-time penalty.
@@
    I classify the top 50 % of occupations with the largest part-time penalties as nonlinear, and the remainder as linear [^9] Nearly all occupations classified as nonlinear exhibit positive part-time penalties, whereas linear occupations typically have negligible or negative penalties..
For example, IT engineers, researchers, pharmaceutical sales representatives, and banking sales representatives are classified as nonlinear occupations. In contrast, construction workers, cooks, and character and CG designers are classified as linear occupations.
@@
99footnotetext: The table provides an unconditional proportion of workers in each category. For example, between the ages of 25 and 59, 4.743% of individuals do not work, while 95.26% are employed. 47.27% of agents work in nonlinear occupations, and 0.2373% of individuals work less than 1,000 hours per year in nonlinear occupations .
@@ 3 Model
In this section, I present the details of the model. I integrate the occupational choice model developed by Jang and Yum [2022] into the retirement decision framework to explain the heterogeneity in the proportion of older workers among overall employees across different occupations.
@@ 3.1 Demographics
A continuum of males is born each year at age 25 (t=1) and lives until age 104 (t=80). A fraction of these individuals begins receiving pension benefits at age 60 (t=36), while the remainder start at age 65 (t=41). Pension eligibility is determined at age 60 as a random shock, remaining unknown until that point. All agents retire from the labor force by age 80 (t=56). Each agent faces a survival shock in every period. Upon death, their bequests are evenly distributed among the remaining survivors.
@@ 3.2 Preferences
Each agent has preferences over consumption, labor supply, and assets which are denoted by $c$, $h$., and $a$, respectively. They decide whether to work and, if working, select occupations. Utility is derived from consumption, while disutility arises from labor supply, which consists of two components: a fixed cost of working, $\xi$, and labor disutility, $\Phi$, which starts to increase at $R+1$. All agents have this utility function:
$u(c,h;\phi,t,a_{0})=\frac{c^{1-\sigma}}{1-\sigma}-\Phi(t)\frac{h^{1+\frac{1}{\gamma}}}{1+\frac{1}{\gamma}}-\xi\mathbbm{1}\{h>0\}$
@@
Agents also have a bequest motive, and all bequests are equally distributed among surviving agents. This model focuses on accidental bequests [^11] I refer to De Nardi [2004] to formulate this bequest motive. and excludes the inheritance of earnings ability and inter-vivo transfers. The utility derived from leaving a bequest is modeled as:
@@ 3.3 Pension
Pension benefits are composed of two terms: the national pension, $\underline{b}$, which is distributed equally to all agents, and employees’ pension insurance, which is based on the agent’s past earnings. The mean of the agent’s past labor earnings is updated each period using the following equation,subject to an upper bound on labor earnings, $\bar{e}$, when calculating pension benefits. Until age 70, $e$ is updated according to the rules of employees’ pension insurance.
@@ 3.4 Efficiency Labor
First, the worker selects an occupation based on the occupation-specific productivity, $m(\eta_{1})$ [^12] NL and L refer to nonlinear and linear occupations, respectively.. A worker draws $\eta_{1}$, an idiosyncratic value, at birth, which remains constant throughout their lifetime. If a worker is well-suited for occupation $j$, they typically remain in the same job until retirement. The occupation-specific productivity is defined as:
@@
Moreover, productivity also depends partly on experience within the occupation. A worker accumulates one unit of experience for each period of work [^13] The process of accumulating experience does not require full-time work, as only 8.9% of males between 25 and 59 work less than 35 hours per week in the data. This full-time work constraint would be necessary if the focus were on analyzing the gender wage gap..
@@
Experience is updated according to the rule: $x=x_{-1}+1$ if $j=j_{-1}$ and $x=0$ if $j\neq j_{-1}$, where $x_{-1}$ represents prior experience. Each unit of an experience adds $\Omega_{j}$ units of an experience premium until reaching the upper bound, $\bar{x}_{j}$ [^14] From the data, I set $x_{NL}=35$ and $x_{L}=29$, which, respectively, denote 59 and 53 years old., which varies across occupations. As long as the worker remains in the same occupation, they continue to accumulate experience. However, if the worker switches to a different occupation, their experience resets, starting from $x=0$.
@@
$f_{j}(t)=\exp\Big({-\pi_{j}(\min\{t,\bar{R}_{j}\}-R)\mathcal{I}_{t>R}}\Big{missing})$
@@ 3.5 Household Problem
Using these features, I construct a household problem that accounts for both the extensive and intensive margins of labor supply, allowing workers to choose their occupations. Productivity increases with age until experience reaches $\bar{x}_{j}$ after which it begins to decline at age $R+1$. Pension eligibility starts unexpectedly for some agents at age 60, while others begin receiving benefits at age 65, with amount based on past earnings. All agents retire by age 80 and live up to 104, facing survival shocks.
@@
$N(a,\phi,\eta_{1},t,e_{-1},p)=max_{c,a\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V^{Y}(a^{\prime},0,0,\phi,\eta_{1},t+1,e,p^{\prime})\Big{]}\bigg{\}}$
@@
$V^{O}(a,t,e_{-1},p^{\prime})=max_{c,a\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V_{O}(a^{\prime},t+1,e,p^{\prime})\Big{]}\bigg{\}}$

### v2

@@
(June 25, 2025)
Older male workers exhibit diverse retirement behaviors across occupations and respond differently to policy changes, influenced significantly by the part-time penalty—wage reduction faced by part-time workers compared to their full-time counterparts. Many older individuals reduce their working hours, and in occupations with high part-time penalties, they tend to retire earlier, as observed in data from Japan and the United States. This study develops a general equilibrium model that incorporates occupational choices, endogenous labor supply, highlighting that the impact on the retirement decision is amplified by the presence of assets and pensions. Using the Japanese Panel Study of Employment Dynamics, I find that cutting employees’ pension benefits reduce aggregate labor supply in occupations with high part-time penalties, reducing overall welfare across the economy. Furthermore, a commonly used policy measure—extending the pension eligibility age—is also found to decrease both output and welfare. In contrast, this paper suggests that increasing income tax credits and exempting pension benefits from income taxation can boost labor supply across all occupations. These policies enhance welfare by raising disposable income relative to the reservation wage.
@@ 1 Introduction
Retirement timings vary significantly across occupations. While some professions tend to see workers retiring at sixty, others have mean retirement ages extending beyond seventy, as shown in Figure 1. Despite this striking heterogeneity, the factors driving such diverse retirement decisions remain unclear. Understanding these underlying factors is a requisite step to examine the impact of policy measures aimed at increasing the labor supply of older workers across the occupational breakdown.
@@
This issue is particularly pressing as labor force is projected to decline further: working-age population in OECD countries is expected to shrink by 11% by 2062 compared to 2022 OECD [2023]. Given the widespread nature of population aging and its uneven impact on labor supply across occupations, it is plausible that policymakers will need to design retirement policies tailored to specific occupational groups. In this context, I investigate how each occupation responds to changes in retirement policy, including an increase in the pension eligibility age and a reduction in pension benefits.
This paper is the first to propose a framework to analyze how different occupations respond to policy changes, highlighting part-time penalties—wage reductions experienced by part-time workers compared to their full-time counterparts—as a key source of this heterogeneity. As illustrated in Figure 3, there is a negative correlation between part-time penalties and share of old workers: a proportion of male workers aged 60 and over among those aged 40 and over. The figure presents a binscatter summarizing occupational trends. The denominator includes both middle-aged and older workers to mitigate the influence of trends in occupational choice. Occupations with smaller part-time penalties tend to have higher rates of old workers as workers face less significant wage reductions when reducing working hours to spend more leisure time as they get older. In other words, in such occupations, workers are likely to retire earlier than in other occupations. This mechanism plays a crucial role in explaining the divergence in retirement timings across occupations.
@@
To classify occupations and compute moments for quantitative analysis, this study utilizes the Japanese Panel Study of Employment Dynamics (JPSED). The JPSED covers more than 200 occupations and provides detailed personal information on each worker, including birth year, sex, education, work history, family status, and more. These rich variables enable precise regressions for classifying occupations. Additionally, the Japanese Household Panel Survey (JHPS/KHPS) supplements the analysis with asset data, which is not available in the JPSED.These datasets allow the study to focus on one of the most rapidly aging populations and to derive policy prescriptions that may be informative for other countries facing similar demographic shifts in the near future. To ensure that similar retirement behaviors are observed in other countries, the IPUMS-CPS (Flood et al. [2024]), which provides data for the United States, is also employed.
@@
I propose several unconventional policies—such as increasing tax credits and exempting pension benefits from income taxation—which are effective in boosting labor supply across both types of occupations, thereby raising output and improving welfare. Notably, these policies reduce tax revenue by less than 3% in general equilibrium.
@@
Figure 3: Binscatter of Part-time Penalty and Share of Old Workers: Males, 2015-2019
@@ 2.1 Data
The Japanese Panel Study of Employment Dynamics (JPSED), compiled by the Recruit Works Institute and released by the University of Tokyo, provides the primary data for occupational classification and model calibration. Spanning the years 2015 to 2022, the dataset includes over 100 individual attributes and approximately 50,000 observations per year, allowing for granular analysis across a range of personal and occupational characteristics.
For occupational classification, the analysis includes all valid observations of individuals aged 25 to 79 from 2015 to 2022, regardless of sex, resulting in a total sample size of 237,897. Older individuals and females are included in this regression to ensure a sufficient number of part-time workers. Among males under age 60, the vast majority are full-time workers, making it difficult to estimate the part-time penalty. To address heterogeneity, the regression controls for sex, age, survey year, marital status, presence of children, and other personal characteristics.
For model calibration, the sample is restricted to males aged 25 to 104 from 2015 to 2019 to avoid the heterogeneous impact of the COVID-19 pandemic on retirement behavior, resulting in a final sample of 93,297 observations.
@@
For robustness checks, IPUMS-CPS data from 2009 to 2024 is used to examine whether a similar pattern is observed in the U.S. The dataset, which includes approximately 130,000 to 220,000 individuals, provides detailed information on personal attributes and work-related characteristics. Occupations are classified based on this data, allowing for a comprehensive analysis of labor market trends and retirement behavior. The data is used as cross-country, and the sample size is 170,071, and the number of occupations is 173.
@@ 2.2 Occupational Classification Strategy
    For each of the more than 200 occupations, I estimate the relationship between hourly wages and weekly working hours using the following regression model. Specifically, I regress hourly wages on a quartic polynomial of weekly working hours, controlling for individual characteristics and fixed effects. Occupations with fewer than 200 observations are excluded from the analysis. After limiting the sample, the number of occupations total to 135.
    $y_{i,j}=\beta_{0,j}+\beta_{1,j}h_{i,j}+\beta_{2,j}h_{i,j}^{2}+\beta_{3,j}h_{i,j}^{3}+\gamma_{j}X_{i,j}+\epsilon_{i,j}$
    Here, $y_{i,j}$ denotes the hourly wage of individual $i$ in occupation $j$, and $h_{i,j}$ represents weekly working hours. The vector $X_{i,j}$ includes control variables such as age (as a polynomial), sex, education, marital status, child status, residential area, and time-fixed effects. The coefficients $\beta_{j}=(\beta_{0,j},\beta_{1,j},\beta_{2,j},\beta_{3,j})$ are estimated separately for each occupation $j$, allowing for occupation-specific wage-hour relationships.
@@
    I calculate the residualized hourly wage difference between individuals working 10 hours per week and those working 40 hours per week. I define this wage difference standarized by the average hourly wage of males aged 25-79 as a part-time penalty for occupational classification. Part-time penalty in occupation j is calculated as follows:
[…]
@@
    I classify the top 50 % of occupations with the largest part-time penalties as nonlinear, and the remainder as linear. Notably, nearly all occupations classified as nonlinear exhibit positive part-time penalties, whereas linear occupations typically show negligible or negative penalties.
For example, IT engineers, researchers, pharmaceutical sales representatives, and banking sales representatives are classified as nonlinear occupations. In contrast, construction workers, cooks, and character and CG designers are classified as linear occupations. A detailed table showing the mapping between each occupation and its nonlinear/linear classification is provided in Appendix A.7. Counterfactual Experiment(Unconventional Policy).
A finer occupational classification is used here, as broader categories encompass a wide variety of jobs with differing part-time penalties. For example, the nature of sales work varies considerably across industries: pharmaceutical sales is classified as nonlinear, whereas insurance sales is linear.
@@
The same phenomenon can be observed in the United States, as shown in Appendix 2. Thanks to greater data availability, the U.S. sample begins in 2009, allowing for a larger number of observations than in Japan, where the data starts in 2015. However, the mechanism is less apparent in the U.S. than in Japan. This is likely because workers in linear occupations in the U.S. may exit the labor force for reasons other than part-time penalties, such as health problems, which appear to play a more significant role than in Japan.
As shown in Figure 16, the labor force participation rate among working-age males in the U.S. is around 90% but begins to decline after age 50. In contrast, the rate in Japan is approximately 95% and remains high until around age 60. At age 60, the labor force participation rate is 71.42% in the U.S., compared to 89.16% in Japan. This discrepancy may reflect differences in health conditions and access to healthcare. These factors are likely to be more prevalent among workers in linear occupations, which primarily consist of low-skilled jobs. In the U.S., such workers may be less able to afford medical expenses and are more likely to experience adverse health conditions, potentially leading to earlier labor force exit in these roles.
@@
88footnotetext: The table provides an unconditional proportion of workers in each category. For example, between the ages of 25 and 59, 4.743% of individuals do not work, while 95.26% are employed. 47.27% of agents work in nonlinear occupations, and 0.2373% of individuals work less than 1,000 hours per year in nonlinear occupations .
@@ 3 Model
This section presents the details of the model. The occupational choice model developed by Jang and Yum [2022] is integrated into a retirement decision framework to account for the heterogeneity in the proportion of older workers across occupations. A distinctive feature of this model is its ability to capture both the extensive and intensive margins of labor supply, while also modeling intergenerational competition for occupational positions— aspects often overlooked in the existing literature, which tends to focus either solely on labor force participation or on individuals around age sixty. These components are important when we consider labor supply of old individuals. The latter part of this section formally defines the concept of a stationary equilibrium.
@@ 3.1 Demographics
t denotes age. A continuum of males is born each year at age 25 (t=1) and lives until age 104 (t=80). A fraction of these individuals begins receiving pension benefits at age 60 (t=36), while the remainder start at age 65 (t=41). Pension eligibility is determined at age 60 as a random shock, remaining unknown until that point. All agents retire from the labor force by age 80 (t=56). Each agent faces a survival shock in every period. Upon death, their bequests are evenly distributed among the remaining survivors.
@@ 3.2 Preferences
Each agent has preferences over consumption and labor supply, which are denoted by $c$ and $h$ respectively. The utility functin is conditional on age and an idiosynacratic intercept of labor disutility function, $\phi$. This paper adopts a separable utility function, following Fan et al. [2022], in contrast to the nonseparable utility specification used in French [2005] and French and Jones [2011]. While those studies focus primarily on individuals near retirement age or within the working-age population, this paper extends the analysis to a broader age range—individuals aged 25 to 104. Even after retirement, individuals continue to spend on healthcare and long-term care, despite a potential sharp decline in overall consumption. A separable utility function is therefore well-suited to capture these life-cycle patterns and better reflect the agents’ choices of consumtption and hours worked. They decide whether to work and, if working, select occupations. Utility is derived from consumption, while disutility arises from labor supply, which consists of two components: a fixed cost of working, $\xi$, and labor disutility, $\Phi$, which starts to increase at $R+1$. All agents have this utility function:
$u(c,h;\phi,t)=\frac{c^{1-\sigma}}{1-\sigma}-\Phi(t)\frac{h^{1+\frac{1}{\gamma}}}{1+\frac{1}{\gamma}}-\xi\mathbbm{1}\{h>0\}$
@@
Agents also have a bequest motive, and all bequests are equally distributed among surviving agents. This is a key driver of saving behavior among older individuals. This model focuses on accidental bequests [^10] I refer to De Nardi [2004] to formulate this bequest motive. and excludes the inheritance of earnings ability and inter-vivo transfers. The utility derived from leaving a bequest is modeled as:
@@ 3.3 Pension
Social security also has a significant impact on retirement decisions. Pension benefits are composed of two terms: the national pension, $\underline{b}$, which is distributed equally to all agents, and employees’ pension insurance, which is based on the agent’s past earnings. The mean of the agent’s past labor earnings, $e$, is updated each period using the following equation,subject to an upper bound on labor earnings, $\bar{e}$, when calculating pension benefits. Until age 70, $e$ is updated according to the rules of employees’ pension insurance.
@@ 3.4 Efficiency Labor
First, the worker selects an occupation based on the occupation-specific productivity, $m(\eta_{1})$. A worker draws $\eta_{1}$, an idiosyncratic value, at birth, which remains constant throughout their lifetime. If a worker is well-suited for occupation $j$, they typically remain in the same job until retirement. The occupation-specific productivity is defined as:
@@
Moreover, productivity also depends partly on experience within the occupation. A worker accumulates one unit of experience for each period of work [^11] The process of accumulating experience does not require full-time work, as only 8.9% of males between 25 and 59 work less than 35 hours per week in the data. This full-time work constraint would be necessary if the focus were on analyzing the gender wage gap.. The experience premium depends on the worker’s previous occupation, denoted by $j_{-1}$.
@@
Experience is updated according to the rule: $x=x_{-1}+1$ if $j=j_{-1}$ and $x=0$ if $j\neq j_{-1}$, where $x_{-1}$ represents prior experience. In other words, if the worker remains in the same occupation, they accumulate one additional unit of experience; if they switch to a different occupation, their occupational experience resets to zero. Each unit of an experience adds $\Omega_{j}$ units of an experience premium until reaching the upper bound, $\bar{x}_{j}$ [^12] Based on the data, I set $x_{NL}=35$ and $x_{L}=29$, which correspond to ages 59 and 53, respectively. These values represent the experience levels at which hourly wages peak in nonlinear and linear occupations. The peak wage ages are interpreted as the maximum effective experience in each occupations., which varies across occupations. As long as the worker remains in the same occupation, they continue to accumulate experience. However, if the worker switches to a different occupation, their experience resets, starting from $x=0$.
@@
$f_{j}(t)=\exp\bigg({-\pi_{j}\Big{(}\min\{t,\bar{R}_{j}\}-R\Big{)}\mathcal{I}_{t>R}}\bigg{missing})$
@@ 3.5 Household Problem
Using these features, I construct a household problem that accounts for both the extensive and intensive margins of labor, allowing workers to choose their occupations. Productivity increases with age until experience reaches $\bar{x}_{j}$ after which it begins to decline at age $R+1$. Pension eligibility begins unexpectedly at age 60 for some agents, while others start receiving benefits at age 65, with benefit levels determined by their historical earnings. This stochastic timing of pension eligibility serves to better align the model with observed data, capturing the gradual exit of older workers from the labor force. All agents are assumed to retire by age 80 and survive up to age 104, facing age-dependent survival risks.
@@
$N(a,\phi,\eta_{1},t,e_{-1},p)=max_{c,a^{\prime}\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V^{Y}(a^{\prime},0,0,\phi,\eta_{1},t+1,e,p^{\prime})\Big{]}\bigg{\}}$
@@
$V^{O}(a,t,e_{-1},p^{\prime})=max_{c,a^{\prime}\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V_{O}(a^{\prime},t+1,e,p^{\prime})\Big{]}\bigg{\}}$